# A Closer Look at Machine Unlearning for Large Language Models

### v3

@@ 2.2 Evaluation metrics
measures the semantic similarity of the model’s output before and after unlearning. Inspired by the semantic textual similarity task (Cer et al., 2017), which aims to assess the degree to which two sentences are semantically equivalent to each other. We use the Sentence-BERT (Reimers & Gurevych, 2019) to get sentence embeddings of output before and after unlearning. Then we calculate their cosine similarity and truncate the value less than 00, denote as $\max(\text{Cos}(g(x;\theta),g(x;\theta_{u}))),0)$. As shown in the second row of Table 1, the unlearned model may add a bunch of unexpected or made-up content after answering the question from the retain set, resulting in a lower CS, though it may have a high ROUGE.
@@
measures the factual correctness of the model’s output for a set of questions relative to the ground truth answers. Text entailment, also known as Natural Language Inference (NLI), is a fundamental task that aims to determine the directional relationship between text fragments. It also plays a crucial role in NLP evaluation (Ferrández et al., 2008; Yao & Barbosa, 2024; Poliak, 2020). Formally, “$t$ entails $h$”($t\Rightarrow h$) if, typically, a human reading text $t$ would infer that the hypothesis $h$ is most likely true (Wikipedia contributors, 2024). Following Liu et al. (2024c), we use a pre-trained NLI model (Sileo, 2023) to predict the relationship between the pair of the model output and the corresponding ground truth for each question. As shown in the third row of Table 1, the wrong name in the output lead to an incorrect fact, its prediction label will be “contradiction”. Then we calculate the proportion of pairs in the set that are predicted to be “entailment” and use this as the entailment score, which should be higher on the retain set and lower on the forget set.
@@ 2.3 Baseline unlearning methods
    $\mathcal{L}_{\text{NPO}}(\mathcal{D}_{\text{F}};\theta)=-\frac{2}{\beta}\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{R}}}\left[\log\sigma\left(-\beta\log\frac{p(y|x;\theta)}{p(y|x;\theta_{\text{ref}})}\right)\right],$ (2)
@@
    $\mathcal{L}_{\text{GD}}(\mathcal{D}_{\text{R}};\theta)$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{F}}}\left[-\log p(y|x;\theta)\right].$ (4)
@@
    $\mathcal{L}_{\text{KL}}(\mathcal{D}_{\text{R}};\theta)$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{F}}}\left[\text{KL}(p(y|x;\theta)\|p(y|x;\theta_{\text{ref}}))\right].$ (5)
@@ 4.1 Lack regularization against rejection templates
Different impacts on the retain set of targeted unlearning. Since the distributions of the forget set and the retain set are similar, i.e., $(\mathcal{X}_{\text{F}},\mathcal{Y}_{\text{F}})\approx(\mathcal{X}_{\text{R}},\mathcal{Y}_{\text{R}})$, decreasing the probability of answers in the forget set, i.e., $\Pr(\mathcal{X}_{\text{F}}|\mathcal{Y}_{\text{F}})$, during untargeted unlearning, will easily decrease $\Pr(\mathcal{X}_{\text{R}}|\mathcal{Y}_{\text{R}})$, as shown in Figure 3. Previous regularization mainly focus on maintaining $\Pr(\mathcal{X}_{\text{R}}|\mathcal{Y}_{\text{R}})$ during the unlearning process. However, we observe that targeted unlearning methods, such as IDK+RT, have little effect on the probability of the original answers, as the primary objective is to increase $\Pr(\mathcal{Y}_{\text{IDK}}|\mathcal{X}_{\text{F}})$, where $\mathcal{Y}_{\text{IDK}}$ is the distribution of rejection templates. The objective of targeted unlearning may also lead to an increase in $\Pr(\mathcal{Y}_{\text{IDK}}|\mathcal{X}_{\text{R}})$, given that $\mathcal{X}_{\text{R}}\approx\mathcal{X}_{\text{F}}$. As shown in Figure 3, the downward trend of ROUGE scores on the retain set is almost consistent with that on the forget set, i.e., output the rejection templates for most questions. Consequently, a new regularization loss needs to be specifically designed for targeted unlearning.

### v5

@@ 2.2 Evaluation metrics
measures the semantic similarity of the model’s output before and after unlearning. Inspired by the semantic textual similarity task (Cer et al., 2017), which aims to assess the degree to which two sentences are semantically equivalent to each other. We use the Sentence-BERT (Reimers & Gurevych, 2019) to get sentence embeddings of output before and after unlearning. Then we calculate their cosine similarity and truncate the value less than 0, denote as $\max(\text{Cos}(g(x;\theta),g(x;\theta_{u}))),0)$. As shown in the second row of Table 1, the unlearned model may add a bunch of unexpected or made-up content after answering the question from the retain set, resulting in a lower CS, though it may have a high ROUGE.
@@
measures the factual correctness of the model’s output for a set of questions relative to the ground truth answers. Text entailment, also known as Natural Language Inference (NLI), is a fundamental task that aims to determine the directional relationship between text fragments. It also plays a crucial role in NLP evaluation (Ferrández et al., 2008; Yao & Barbosa, 2024; Poliak, 2020). Formally, “$t$ entails hh”($t\Rightarrow h$) if, typically, a human reading text $t$ would infer that the hypothesis $h$ is most likely true (Wikipedia contributors, 2024). Following Liu et al. (2024c), we use a pre-trained NLI model (Sileo, 2023) to predict the relationship between the pair of the model output and the corresponding ground truth for each question. As shown in the third row of Table 1, the wrong name in the output lead to an incorrect fact, its prediction label will be “contradiction”. Then we calculate the proportion of pairs in the set that are predicted to be “entailment” and use this as the entailment score, which should be higher on the retain set and lower on the forget set.
@@ 2.3 Baseline unlearning methods
    $\mathcal{L}_{\text{NPO}}(\mathcal{D}_{\text{F}};\theta)=-\frac{2}{\beta}\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{F}}}\left[\log\sigma\left(-\beta\log\frac{p(y|x;\theta)}{p(y|x;\theta_{\text{ref}})}\right)\right],$ (2)
@@
    $\mathcal{L}_{\text{GD}}(\mathcal{D}_{\text{R}};\theta)$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{R}}}\left[-\log p(y|x;\theta)\right].$ (4)
@@
    $\mathcal{L}_{\text{KL}}(\mathcal{D}_{\text{R}};\theta)$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{R}}}\left[\text{KL}(p(y|x;\theta)\|p(y|x;\theta_{\text{ref}}))\right].$ (5)
@@ 4.1 Lack regularization against rejection templates
Different impacts on the retain set of targeted unlearning. Since the distributions of the forget set and the retain set are similar, i.e., $(\mathcal{X}_{\text{F}},\mathcal{Y}_{\text{F}})\approx(\mathcal{X}_{\text{R}},\mathcal{Y}_{\text{R}})$, decreasing the probability of answers in the forget set, i.e., $\Pr(\mathcal{Y}_{\text{F}}|\mathcal{X}_{\text{F}})$, during untargeted unlearning, will easily decrease $\Pr(\mathcal{Y}_{\text{R}}|\mathcal{X}_{\text{R}})$, as shown in Figure 3. Previous regularization mainly focus on maintaining $\Pr(\mathcal{Y}_{\text{R}}|\mathcal{X}_{\text{R}})$ during the unlearning process. However, we observe that targeted unlearning methods, such as IDK+RT, have little effect on the probability of the original answers, as the primary objective is to increase $\Pr(\mathcal{Y}_{\text{IDK}}|\mathcal{X}_{\text{F}})$, where $\mathcal{Y}_{\text{IDK}}$ is the distribution of rejection templates. The objective of targeted unlearning may also lead to an increase in $\Pr(\mathcal{Y}_{\text{IDK}}|\mathcal{X}_{\text{R}})$, given that $\mathcal{X}_{\text{R}}\approx\mathcal{X}_{\text{F}}$. As shown in Figure 3, the downward trend of ROUGE scores on the retain set is almost consistent with that on the forget set, i.e., output the rejection templates for most questions. Consequently, a new regularization loss needs to be specifically designed for targeted unlearning.